26. A table tennis ball with mass m = 50 gr is left free from a height of h = 2 m. During the perpendicular stroke, lasting delta t = 0.02s, the velocity changes direction but not the value. a) Calculate the impact forces. b) Argue by calculations, if the effect of the gravitational force on the change of impulse, during the shock, can not be taken into account.


1
Expert's answer
2021-02-02T17:05:37-0500

a)


"F=0.05\frac{\sqrt{2(9.8)(2)}}{0.02}=15.65\ N"

b)


"W=0.05(9.8)=0.49\ N"